OpenStudy (anonymous):

Use factoring and the zero product property to solve equation. 2x^2 -9x -18 =0

OpenStudy (xxdragoonxx):

2x+3=0 2x+3-3=0-3 2x=-3 2x/2=-3/2 x = -3/2 x - 6 = 0 x - 6 + 6 = 0 + 6 x = 6 x = 6, x = -3/2